Sodium and chlorine combine to form sodium chloride. How many grams of sodium chloride are produced if 8.3 moles of chlorine react?
1
Expert's answer
2020-04-02T09:04:15-0400

n(Cl2) = 8.3 moles


Solution:

Sodium + chlorine = sodium chloride.

The equation of a chemical reaction:

2Na + Cl2 = 2NaCl

According to the chemical equation: n(Na)/2 = n(Cl2) = n(NaCl)/2.

Therefore,

n(NaCl) = 2*n(Cl2) = 2 * (8.3 moles) = 16.6 moles.

moles of NaCl = n(NaCl) = m(NaCl) / M(NaCl).

M(NaCl) = Ar(Na) + Ar(Cl) = 23 + 35.5 = 58.5 (g/mol).

Then,

m(NaCl) = n(NaCl) * M(NaCl) = (16.6 moles) * (58.5 g/mol) = 971.1 g

m(NaCl) = 971.1 g.


Answer: 971.1 of sodium chloride (NaCl) are produced.